VPTOVF(1) VPTOVF(1) NAME vptovf-convertvirtualpropertyliststovirtualfontmetrics SYNOPSIS vptovf[-verbose]vplfilenamevfmfilenametfmfilename DESCRIPTION Thevptovfprogramtranslatesa(human-oriented)propertylistfiletoapairof(program-oriented) filesintheVF(virtualfont)andTFM(TeXfontmetric)formats.Thus,avirtualfontfilecanbe editedanditsexactcontentscanbedisplayedmnemonically.Newvirtualfonts,whichmapcharactersasseenbyTeXintoanarbitrarysequenceoflow-leveltypesettingoperations,canalsobe createdinthisway. Allthreefilenames,vplfilename,vffilename,andtfmfilename,mustbecomplete.Noadding ofdefaultextensionsorpathsearchingisdone. OPTIONS Withoutthe-verboseoption,vptovfoperatessilently.Withit,abannerandprogressreportare printedonstdout.Seetex(1)forthedetailsofspecifyingoptions. SEEALSO vftovp(1) DonaldKnuth,VirtualFonts:MoreFunforGrandWizards,TUGboat11(1),April1990. AUTHORS DonaldKnuthwrotetheprogram,basedinpartonanideaofDavidFuchs,startingwiththecode forpltotf(1).KarlBerryadapteditforcompilationwithweb2c. 7Jan92 1